# LEKA: LLM-Enhanced Knowledge Augmentation

### v2

@@ Abstract
Humans excel in analogical learning and knowledge transfer and, more importantly, possess a unique understanding of identifying appropriate sources of knowledge. From a model’s perspective, this presents an interesting challenge. If models could autonomously retrieve knowledge useful for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge. However, filling models with knowledge is relatively straightforward—it simply requires more training and accessible knowledge bases. The more complex task is teaching models about which knowledge can be analogized and transferred. Therefore, we design a knowledge augmentation method, LEKA, for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge. This LEKA method extracts key information from the target domain’s textual information, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures. We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in reducing computational costs, automating data alignment, and optimizing transfer learning outcomes.
@@ 1 Introduction
Humans are good at identifying proper sources of knowledge. This is an ability that is rooted in our capability for analogical reasoning and knowledge management. In contrast, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information. This gap presents a challenge in enriching domain knowledge and enhancing data augmentation.
Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures Tang et al. (2020). In knowledge augmentation, knowledge transfer is an important method improving learning performance by transferring knowledge from external information sources Khodaee et al. (2024). In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness using relevant external information. In these scenarios, knowledge augmentation, especially knowledge transfer, can reduce reliance on extensive target domain data by strategically selecting sources matching target needs.
@@
Figure 1: Example of LEKA. We adopt an LLM to retrieve proper source domain data to transfer knowledge to a data-limited target domain. The LLM extracts the key information of target data to retrieve for a relevant dataset; then we adopt the LLM for harmonization.
@@
To address the challenges above, we design LLM-Enhanced Knowledge Augmentation (LEKA), a novel and automated data retrieval and augmentation method by knowledge augmentation. Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset-RAG in an external database to efficiently extract knowledge relevant to the target domain dataset. The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain on feature space and marginal probability measures to enhance downstream task learning performance. The data harmonization reduces the structural and semantic differences between the source and target data. This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
As shown in Figure 1, consider a target domain focused on predicting a rare cancer, direct learning is challenging due to insufficient sample information. To augment the data with external information, would typically require human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets. In contrast, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information. This retrieved dataset possesses a more complete and transparent feature space with a structure similar to the target domain. The LLM then adjusts the retrieved data based on the characteristics of the target domain. In this way, we achieve a highly harmonized source dataset for knowledge augmentation.
@@ 2.1 Knowledge Transfer
Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task Alyafeai et al. (2020); Wang et al. (2022, 2025). Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences. Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge He et al. (2021); Han et al. (2021); Yordanov et al. (2021). Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the model’s effectiveness by providing a relevant starting point. While many methods exist to adjust existing knowledge within the same or across different domains to improve transfer learning outcomes, the process of retrieving certain knowledge still heavily relies on manual effort.
@@ 2.2 Data Harmonizing with LLMs
Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency Cao et al. (2009); Zhang et al. (2024a); Liu et al. (2021); Durante et al. (2024). However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process Feng et al. (2021); Xie et al. (2024); Zhang et al. (2025a). Furthermore, the risk of overfitting remains a concern, as models may become overly specialized in training data nuances, reducing their effectiveness on new datasets.
@@ 2.3 Retrieval Augmented Generation for Knowledge Augmentation
Retrieval Augmented Generation (RAG) Lewis et al. (2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation Hu and Lu (2024). This technique aids LLMs in tasks that demand deep and specific domain knowledge Zhang et al. (2024b); Huang and Huang (2024). RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains Siriwardhana et al. (2023). RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized. In this way it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
@@ 4 Methodology
Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of the target dataset, then 2) retrieves datasets in libraries and 3) processes data harmonization. With harmonized datasets, we can transfer knowledge from the source dataset we construct to enhance learning on the target dataset.
@@ 4.1 Dataset Retrieval
In this phase, our primary objective is to identify and retrieve a source dataset $\mathcal{D}_{S}$ similar to the target dataset $\mathcal{D}_{T}$ for effective pre-training in knowledge transfer scenarios. To achieve this, we utilize an LLM $\mathcal{A}$ to analyze and retrieve data that matches the structural properties and purposes of $\mathcal{D}_{T}$. A straightforward method is to embed the target dataset $\mathcal{D}_{T}$ by $\mathcal{A}$ to capture its essential features. However, this approach can be computationally intensive and prone to errors. Inspired by the strategies outlined in Zhang et al. (2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$. These textual elements provide rich insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focuses on the semantic content of $\mathcal{D}_{T}$.
@@ 4.2 Data Harmonization
Now that we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM. This alignment process is to adjust the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$. This process involves transformations of features and adjustments of distributions. Here we take $\mathcal{D}_{k}$ as an example.
@@
Here, $\kappa(\cdot,\cdot)$ is a kernel function, $x_{i}^{k}$ and $x_{j}^{k}$ are the $i$-th and $j$-th feature vectors in the source dataset $\mathcal{D}_{k}$. We adopt a Gaussian kernel $\kappa(x,y)=\exp(-\gamma\|x-y\|^{2})$, with $\gamma$ being a positive bandwidth parameter. This kernel function can effectively measure the similarity between pairs of points in high-dimensional spaces and captures both linear and non-linear relationships. Thus, it is capable of handling the complexities inherent in high-dimensional data. Specifically, compared with Euclidean distances, kernel distances can capture the geometric structure of the data manifold. It provides a more robust and informative measure of similarity for the source-target alignment.
@@
Our LLM then aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets. The LLM analyzes and refines the textual information, including labels, feature names, and classification probability distribution, etc. To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
[…]
@@
Moreover, the Wasserstein distance is particularly useful when dealing with distributions whose support sets (i.e., the effective range or set of the distributions) do not fully overlap. In real-world data applications, it is common for the source and target datasets to originate from different distributions. Their data points do not align perfectly and cover the same areas. In these cases, traditional distance metrics like the Euclidean distance may not effectively reflect the true differences between the two datasets, as they merely measure differences in position but ignore the overall structure of the data distributions.
With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain. Thus, it enhances the overall performance of our transfer learning models. Further, we optimize data handling and boost the adaptability and accuracy of transfer learning for adoption in various applications
@@ 4.3 Knowledge Transfer and Evaluation
After aligning the source dataset $\mathcal{D}_{k}^{\prime}$ with the target dataset $\mathcal{D}_{T}$, and adjusting the marginal probability measures $\mathcal{P}_{k}^{\prime}$ to $\mathcal{P}_{T}$, we proceed to integrate these aligned datasets into the transfer learning model. At this stage, our goal is to effectively transfer knowledge from $\mathcal{D}_{k}^{\prime}$ to $\mathcal{D}_{T}$ by minimizing a defined loss function while adapting the model to the target domain.
@@
During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, a dataset that incorporates domain-specific characteristics to reduce domain shift. We define the optimization process as:
$\theta^{*}$ $=\arg\min_{\theta}\bigg{(}\alpha\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]$ (9)
@@
This formula aims to fine-tune the decision function $f_{T_{j}}$ by minimizing the weighted sum of expected losses across the datasets. The loss function $\mathcal{L}(f_{T_{j}}(x;\theta),y)$ evaluates prediction accuracy, guiding the adjustment of parameters. The expectations $\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}$ and $\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}$ represent the mean losses over the source and target datasets, respectively. The weighting factor $\alpha$ adjusts the relative influence of each dataset and enables flexible adaptation between leveraging established knowledge from the source and integrating new data from the target. In this way, we utilize the transferred knowledge to enhance model performance on the target task.
@@
where $\eta$ is the learning rate. The gradients are computed based on both datasets, which allows the model to learn both from the aligned source data and the target data.
After the parameter optimization and backpropagation processes, we evaluate the effectiveness of the transfer learning process; we adopt a performance metric $\phi$ on the target dataset:
@@
Table 1: Datasets description. Here we use four datasets from medical and economic domains.
@@ 5 Experiments
In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA. First, we compare the performance of the LEKA against several baseline methods on four downstream tasks. Second, we present the correlations between several target domains and their retrieved source domains. Finally, we discuss the reason for the improvement of performance.
@@ 5.1 Experiment Settings
We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) Wolberg et al. (1995), (2) Heart Disease (HD) Janosi et al. (1989), (3) Vehicle Insurance Data (VID) Bhatt (2019), and (4) Telco Customer Churn (TCC) BlastChar (2018). We show the detailed information about features of the datasets in Table 1.
@@
In our setup for data synthesis and model training, we utilize GPT-4o as the query generator, combined with the Exa API to fetch web pages containing datasets from Kaggle and the UCI Machine Learning Repository that may be suitable for knowledge transfer. We extract dataset descriptions from web pages, and GPT-4o assesses their potential for knowledge transfer. Additionally, GPT-4o serves as a generator for executable code, performing up to five code generations. For our models, we configure TN, TT, and FTT with a batch size of $512$ for the VID and TCC datasets and a batch size of $32$ for the BCW dataset, a maximum of $100$ epochs, and employ early stopping with a patience of $20$. The learning rate is set at the default $0.02$ for pytorch_tabnet. For the RF and GBDT models, the number of trees is set to $100$, with GBDT also configured with a learning rate of $0.1$ and a max depth of $3$. TTab is set with a maximum of $50$ epochs, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
Metrics Model GBDT RF TN XB TT

### v3

@@ Abstract
Humans excel in analogical learning and knowledge transfer and, more importantly, possess a unique understanding of identifying appropriate sources of knowledge. From a model’s perspective, this presents a unique challenge. If models could autonomously retrieve knowledge relevant for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge. However, filling models with knowledge is relatively straightforward—it simply requires more training and accessible knowledge bases. The more complex task is teaching models about which knowledge can be analogized and transferred. Therefore, we design a knowledge augmentation method, LEKA, for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge. This LEKA method extracts key information from the target domain’s textual information, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures. We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in automating data alignment and optimizing transfer learning outcomes.
@@ 1 Introduction
Humans are good at identifying relevant sources of knowledge. This is an ability that is rooted in our capability for analogical reasoning and knowledge management. In contrast, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information. This gap presents a challenge in enriching domain knowledge and enhancing data augmentation.
Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures Tang et al. (2020). Knowledge transfer is a crucial method within knowledge augmentation for improving learning performance by transferring knowledge from external information sources Khodaee et al. (2024). In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness using relevant external information. In these scenarios, knowledge augmentation, especially knowledge transfer, can reduce reliance on extensive target domain data by strategically selecting sources matching target needs.
@@
Figure 1: Example of LEKA. We adopt an LLM to retrieve proper source domain data to transfer knowledge to a data-limited target domain. The LLM extracts the key information of the target data to retrieve a relevant dataset; then, we adopt the LLM for harmonization.
@@
To address the challenges above, we design LLM-Enhanced Knowledge Augmentation (LEKA), a novel and automated data retrieval and augmentation method by knowledge augmentation. Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset-RAG in an external database to efficiently extract knowledge relevant to the target domain dataset. The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain in feature space and marginal probability measures to enhance downstream task learning performance. The data harmonization reduces the structural and semantic differences between the source and target data. This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
As shown in Figure 1, consider a target domain focused on predicting a rare cancer, so direct learning is challenging due to insufficient sample information. To augment the data with external information would typically require human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets. In contrast, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information. This retrieved dataset possesses a more complete and transparent feature space with a structure similar to the target domain. The LLM then adjusts the retrieved data based on the characteristics of the target domain. In this way, we achieve a highly harmonized source dataset for knowledge augmentation.
@@ 2.1 Knowledge Transfer
Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task Alyafeai et al. (2020); Wang et al. (2022, 2025). Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences. Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge Han et al. (2021); Yordanov et al. (2021). Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the model’s effectiveness by providing a relevant starting point. While many methods exist to adjust existing knowledge within the same or across different domains to improve transfer learning outcomes, the process of retrieving certain knowledge still heavily relies on manual effort.
@@ 2.2 Data Harmonizing with LLMs
Data harmonization applying LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency Cao et al. (2009); Zhang et al. (2024a); Liu et al. (2021); Durante et al. (2024). However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process Feng et al. (2021); Xie et al. (2024); Zhang et al. (2025a). Furthermore, the risk of overfitting remains a concern, as models may become overly specialized in training data nuances, reducing their effectiveness on new datasets.
@@ 2.3 Retrieval Augmented Generation for Knowledge Augmentation
Retrieval Augmented Generation (RAG) Lewis et al. (2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation Hu and Lu (2024). This technique aids LLMs in tasks that demand deep and specific domain knowledge Zhang et al. (2024b); Huang and Huang (2024). RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains Siriwardhana et al. (2023). RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized. In this way, it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
@@ 4 Methodology
Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of the target dataset, then 2) retrieves datasets in libraries, and 3) processes data harmonization. With harmonized datasets, we can transfer knowledge from the source dataset we construct to enhance learning on the target dataset.
@@ 4.1 Dataset Retrieval
In this phase, our primary objective is to identify and retrieve a source dataset $\mathcal{D}_{S}$ similar to the target dataset $\mathcal{D}_{T}$ for effective pre-training in knowledge transfer scenarios. To achieve this, we utilize an LLM $\mathcal{A}$ to analyze and retrieve data that matches the structural properties and purposes of $\mathcal{D}_{T}$. A straightforward method is to embed the target dataset $\mathcal{D}_{T}$ by $\mathcal{A}$ to capture its essential features. However, this approach can be computationally intensive and prone to errors. Inspired by the strategies outlined in Zhang et al. (2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$. These textual elements provide insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focus on the semantic content of $\mathcal{D}_{T}$.
@@ 4.2 Data Harmonization
Now that we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM. This alignment process adjusts the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$. This process involves transformations of features and adjustments of distributions. Here we take $\mathcal{D}_{k}$ as an example.
@@
Here, $\kappa(\cdot,\cdot)$ is a kernel function, $x_{i}^{k}$ and $x_{j}^{k}$ are the $i$-th and $j$-th feature vectors in the source dataset $\mathcal{D}_{k}$. We adopt a Gaussian kernel $\kappa(x,y)=\exp(-\gamma\|x-y\|^{2})$, with $\gamma$ being a positive bandwidth parameter. This kernel function can effectively measure the similarity between points in high-dimensional spaces and captures both linear and non-linear relationships. Thus, it can handle the complexities inherent in high-dimensional data. Compared with Euclidean distances, kernel distances can capture the geometric structure of the data manifold. It provides a more robust and informative similarity measure for the source-target alignment.
@@
Our LLM aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets. The LLM analyzes and refines textual information such as labels, feature names, and classification probability distributions. To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
$W(\mathcal{P}_{k},\mathcal{P}_{T})=\inf_{\gamma\in\Gamma(\mathcal{P}_{k},\mathcal{P}_{T})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|\,d\gamma(x,y),$ (6)
@@
Moreover, the Wasserstein distance is advantageous when dealing with distributions whose support sets (i.e., the effective range or set of the distributions) do not fully overlap. In real-world data applications, it is common for the source and target datasets to originate from different distributions. Their data points do not align perfectly and cover the same areas. In these cases, traditional distance metrics like the Euclidean distance poorly reflect the actual differences between the two datasets, as they merely measure differences in position but ignore the overall structure of the data distributions.
With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain. Thus, it enhances the overall performance of our transfer learning models. Further, we optimize data handling and boost the adaptability and accuracy of transfer learning for adoption in various applications.
@@ 4.3 Knowledge Transfer and Evaluation
After aligning the source dataset $\mathcal{D}_{k}^{\prime}$ with the target dataset $\mathcal{D}_{T}$, and adjusting the marginal probability measures $\mathcal{P}_{k}^{\prime}$ to $\mathcal{P}_{T}$, we proceed to integrate these aligned datasets into the transfer learning model. At this stage, our goal is to transfer knowledge from $\mathcal{D}_{k}^{\prime}$ to $\mathcal{D}_{T}$ by minimizing a defined loss function while adapting the model to the target domain.
@@
During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, incorporating domain-specific characteristics to reduce domain shift. We define the optimization process as:
$\theta^{*}$ $=\arg\min_{\theta}\bigg(\alpha\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]$
@@
This formula aims to fine-tune the decision function $f_{T_{j}}$ by minimizing the weighted sum of expected losses across the datasets. The loss function $\mathcal{L}(f_{T_{j}}(x;\theta),y)$ evaluates prediction accuracy, guiding the adjustment of parameters. The expectations $\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}$ and $\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}$ represent the mean losses over the source and target datasets, respectively. The weighting factor $\alpha$ adjusts the relative influence of each dataset and enables flexible adaptation between leveraging established knowledge from the source and integrating new data from the target. In this way, we apply the transferred knowledge to enhance model performance on the target task.
@@
where $\eta$ is the learning rate. The gradients are computed based on both datasets, which allows the model to learn from both the aligned source data and the target data.
After the parameter optimization and backpropagation, we evaluate the effectiveness of the transfer learning process; we adopt a performance metric $\phi$ on the target dataset:
@@
Table 1: Datasets description. Here we use four datasets from the medical and economic domains.
@@ 5 Experiments
In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA. First, we compare the LEKA against several baseline methods on four downstream tasks. Second, we present the correlations between several target domains and their retrieved source domains. Finally, we discuss the reason for performance improvement.
@@ 5.1 Experiment Settings
We evaluate our method on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) Wolberg et al. (1995), (2) Heart Disease (HD) Janosi et al. (1989), (3) Vehicle Insurance Data (VID) Bhatt (2019), and (4) Telco Customer Churn (TCC) BlastChar (2018). We show the detailed information about the features of the datasets in Table 1.
@@
In our setup for data synthesis and model training, we utilize GPT-4o OpenAI (2024) as the query generator, combined with the Exa API Exa (2024) to fetch web pages containing datasets from Kaggle Kaggle (2024) and the UCI Machine Learning Repository University of California, Irvine (2024) that may be suitable for knowledge transfer. We extract dataset descriptions from web pages, and GPT-4o assesses their potential for knowledge transfer. Additionally, GPT-4o serves as a generator for executable code, performing up to five code generations. For our models, we configure TN, TT, and FTT with a batch size of $512$ for the VID and TCC datasets and a batch size of $32$ for the BCW dataset, a maximum of $100$ epochs, and employ early stopping with a patience of $20$. The learning rate is set at the default $0.02$ for pytorch_tabnet. For the RF and GBDT models, the number of trees is set to $100$, with GBDT also configured with a learning rate of $0.1$ and a max depth of $3$. TTab is set with a maximum of $50$ epochs, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
Metrics Model GBDT RF TN XB $T$